# From Semantic To Instance: A Semi-Self-Supervised Learning Approach

### v1

@@ 4 Results
SynModel YOLOv9e $\textit{LateStage}_{te}$ GLM 86.5 82.0 88.2 61.2
$\textit{GHD}_{te}$ GLM 95.2 95.3 97.9 77.0
RoAModel SynModel $\textit{LateStage}_{te}$ GLM 93.2 93.3 95.7 80.7
$\textit{GHD}_{te}$ GLM 98.1 96.5 98.5 85.3
PseModel SynModel $\textit{LateStage}_{te}$ GLM 77.1 71.7 77.5 49.0
$\textit{GHD}_{te}$ GLM 95.8 95.2 98.0 76.6
@@
RoAModel was built upon SynModel by being fine-tuned on the rotation-augmented dataset $\textit{Rotated}_{tr}$ and $\textit{Rotated}_{va}$. RoAModel obtained the high-performance of $95.7\%$ mAP@50 on our single-domain $\textit{LateStage}_{te}$ set and a mAP@50 of $98.5\%$ on $\textit{GHD}_{te}$. Figure 5 displays the prediction performance of this model on samples from all $18$ domains of $\textit{GHD}_{te}$. Considering the better mAP@50-95 scores of this model on both sets, Figure 6 visually presents the prediction quality of the RoAModel in comparison to BaseModel on a few samples chosen from $\textit{LateStage}_{te}$ and $\textit{GHD}_{te}$ test sets. Regarding the quality assessment on $\textit{HandHeld}_{qa}$ and $\textit{GHD}_{qa}$oAModel visually demonstrates substantial accuracy across all $28$ data domains (please see Figures 7 and  8). We conducted an additional comparative experiment by training PseModel on SynModel, using $\textit{Pseudo}_{tr}$ and $\textit{Pseudo}_{va}$. This model underperformed RoAModel when evaluated on both test sets.
@@
Table 3: Performance comparison of the same YOLOv9e architectures trained from scratch using RGB (base model) and GLM (our approach) versions of COCO dataset [19].
@@
None GLM 70.6 58.7 68.0 55.4
We additionally experimented with our proposed approach on a general-purpose dataset and task using the Microsoft COCO 2017 dataset [19], extending beyond precision agriculture to asses it’s applicability across broader domains. We trained two models, one using the dataset’s RGB version of training and validation sets and one applying our proposed GLM version. Table 3 presents model performances on the validation set of this dataset, illustrating that our proposed approach achieved improvements in up to $12.6\%$ of mAP@50 and over $17.8\%$ in mAP@50-95. Additionally, Figure 9 illustrates a comparison of prediction performances between the RGB and GLM models, displayed in the second and third rows, respectively, with the ground truth shown in the first row. Each column in the figure represents a distinct sample, annotated with the ground truth labels, annotations generated by the base RGB model, and annotations produced by the GLM model developed using our proposed approach.
Figure 9: Prediction performance of the COCO models on the Microsoft COCO 2017 dataset [18]. Our proposed GLM approach consistently achieved superior segmentation performance (columns B, C, D, and E), and obtained higher confidence scores for detected objects (A through F). In some cases both RGB and GLM models failed to detect objects of interest, examples include the truck in column D and people in F. In addition, near-perfect segmentation performance was observed in cases such as A.
@@ 5 Discussion and Limitations
Beyond the agricultural domain, evaluating our proposed approach on a general-purpose instance segmentation task using the MS COCO dataset [19] supports its applicability across diverse fields. Although only binary masks, rather than semantic masks, were used to generate the GLM version of the COCO dataset for model development, this approach achieved a $17.8\%$ improvement in performance compared to the baseline YOLO model trained on the RGB version under identical training and evaluation conditions for fair comparison. As our approach originally advocated, the generation of GLM images using semantic masks instead of binary masks could further enhance CocoModel performance, suggesting that future studies should explore the use of semantic masks and additional color spaces to create GLM images.

### v2

@@ 4 Results
SynModel YOLOv9e $\textit{LateStage}_{te}$ GLMask 86.5 82.0 88.2 61.2
$\textit{GHD}_{te}$ GLMask 95.2 95.3 97.9 77.0
RoAModel SynModel $\textit{LateStage}_{te}$ GLMask 93.2 93.3 95.7 80.7
$\textit{GHD}_{te}$ GLMask 98.1 96.5 98.5 85.3
PseModel SynModel $\textit{LateStage}_{te}$ GLMask 77.1 71.7 77.5 49.0
$\textit{GHD}_{te}$ GLMask 95.8 95.2 98.0 76.6
@@
RoAModel was built upon SynModel by being fine-tuned on the rotation-augmented dataset $\textit{Rotated}_{tr}$ and $\textit{Rotated}_{va}$. RoAModel obtained the high-performance of $95.7\%$ mAP@50 on our single-domain $\textit{LateStage}_{te}$ set and a mAP@50 of $98.5\%$ on $\textit{GHD}_{te}$. Figure 5 displays the prediction performance of this model on samples from all $18$ domains of $\textit{GHD}_{te}$.
Table 3: Performance evaluation of our BaseModel (base) and RoAModel (best) models across the $18$ diverse domains of the $GHD_{te}$ dataset.
Domain Model Pretrained mAP@50
[Uncaptioned image] BaseModel YOLOv9e 50.5
RoAModel SynModel 96.9
[Uncaptioned image] BaseModel YOLOv9e 57.5
RoAModel SynModel 97.2
[Uncaptioned image] BaseModel YOLOv9e 71.1
RoAModel SynModel 98.0
[Uncaptioned image] BaseModel YOLOv9e 14.9
RoAModel SynModel 99.5
[Uncaptioned image] BaseModel YOLOv9e 28.9
RoAModel SynModel 99.0
[Uncaptioned image] BaseModel YOLOv9e 60.1
RoAModel SynModel 99.3
[Uncaptioned image] BaseModel YOLOv9e 29.9
RoAModel SynModel 99.5
[Uncaptioned image] BaseModel YOLOv9e 35.5
RoAModel SynModel 99.0
[Uncaptioned image] BaseModel YOLOv9e 28.1
RoAModel SynModel 97.7
Domain Model Pretrained mAP@50
[Uncaptioned image] BaseModel YOLOv9e 53.4
RoAModel SynModel 98.7
[Uncaptioned image] BaseModel YOLOv9e 16.4
RoAModel SynModel 98.4
[Uncaptioned image] BaseModel YOLOv9e 50.7
RoAModel SynModel 99.0
[Uncaptioned image] BaseModel YOLOv9e 33.2
RoAModel SynModel 99.3
[Uncaptioned image] BaseModel YOLOv9e 08.7
RoAModel SynModel 99.5
[Uncaptioned image] BaseModel YOLOv9e 24.9
RoAModel SynModel 98.3
[Uncaptioned image] BaseModel YOLOv9e 62.6
RoAModel SynModel 95.8
[Uncaptioned image] BaseModel YOLOv9e 34.4
RoAModel SynModel 98.9
[Uncaptioned image] BaseModel YOLOv9e 18.1
RoAModel SynModel 99.5
To evaluate the generalization capability and robustness of the proposed framework, we conducted a comprehensive per-domain performance analysis across the $18$ diverse acquisition environments in the $\textit{GHD}_{te}$ dataset. As summarized in Table 3, the RoAModel demonstrates significant and consistent performance gains over the RGB BaseModel in every evaluated domain. While the BaseModel exhibits high variance and struggles in several challenging environments (Arvalis_5 and Arvalis_4), our approach maintains high $mAP@50$ scores, effectively bridging the domain gap through the combination of GLMask representation and rotation-augmented adaptation.
Considering the better mAP@50-95 scores of this model on both sets, Figure 6 visually presents the prediction quality of the RoAModel in comparison to BaseModel on a few samples chosen from $\textit{LateStage}_{te}$ and $\textit{GHD}_{te}$ test sets. Regarding the quality assessment on $\textit{HandHeld}_{qa}$ and $\textit{GHD}_{qa}$RoAModel visually demonstrates substantial accuracy across all $28$ data domains (please see Figures 7 and  8). We conducted an additional comparative experiment by training PseModel on SynModel, using $\textit{Pseudo}_{tr}$ and $\textit{Pseudo}_{va}$. This model underperformed RoAModel when evaluated on both test sets.
@@
Table 4: Performance comparison of the same YOLOv9e architectures trained from scratch using RGB (base model) and GLMask (our approach) versions of COCO dataset [19].
@@
None GLMask 70.6 58.7 68.0 55.4
We additionally experimented with our proposed approach on a general-purpose dataset and task using the Microsoft COCO 2017 dataset [19], extending beyond precision agriculture to assess its applicability across broader domains. We trained two models, one using the dataset’s RGB version of training and validation sets and one applying our proposed GLMask version. Table 4 presents model performances on the validation set of this dataset, illustrating that our proposed approach achieved improvements in up to $12.6\%$ of mAP@50 and over $17.8\%$ in mAP@50-95. Additionally, Figure 9 illustrates a comparison of prediction performances between the RGB and GLMask models, displayed in the second and third rows, respectively, with the ground truth shown in the first row. Each column in the figure represents a distinct sample, annotated with the ground truth labels, annotations generated by the base RGB model, and annotations produced by the GLMask model developed using our proposed approach.
Figure 9: Prediction performance of the COCO models on the Microsoft COCO 2017 dataset [18]. Our proposed GLMask approach consistently achieved superior segmentation performance (columns B, C, D, and E), and obtained higher confidence scores for detected objects (A through F). In some cases both RGB and GLMask models failed to detect objects of interest, examples include the truck in column D and people in F. In addition, near-perfect segmentation performance was observed in cases such as A.
@@ 5 Discussion and Limitations
Beyond the agricultural domain, evaluating our proposed approach on a general-purpose instance segmentation task using the MS COCO dataset [19] supports its applicability across diverse fields. Although only binary masks, rather than semantic masks, were used to generate the GLMask version of the COCO dataset for model development, this approach achieved a $17.8\%$ improvement in performance compared to the baseline YOLO model trained on the RGB version under identical training and evaluation conditions for fair comparison. As our approach originally advocated, the generation of GLMask images using semantic masks instead of binary masks could further enhance CocoModel performance, suggesting that future studies should explore the use of semantic masks and additional color spaces to create GLMask images.
@@ 6 Conclusion
Acknowledgements
Author Contributions: Conceptualization and Problem Definition, F. Maleki, L. Jin, I. Stavness; Data collection and annotation, K. Najafian; Model development and experimentation, K. Najafian; Visualization, K.N.; Writing original draft preparation and literature review, K. Najafian; Writing—review and editing, F. Maleki, L. Jin, and I. Stavness; resources, L. Jin and I. Stavness; Supervision, F. Maleki, L. Jin, and I. Stavness. All authors have reviewed and approved the final version of the manuscript.
Conflict of Interest Statement: The authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
Funding: This research is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC), grant number 2019-06424 (LJ), 2019-07017 (IS), and 2024-04966 (FM).
Data Statement: The data and code supporting the findings of this study will be made available upon acceptance of the paper.
Sex and Gender Considerations: This study analyzes agricultural data collected via cameras and drones and does not involve human or animal subjects. As such, sex and gender considerations do not apply to the research. The other dataset used in this study is publicly available and does not require sex or gender-based analysis.
Ethical Considerations: This study does not involve human or animal subjects. It utilizes agricultural data collected through cameras and a widely recognized, publicly available dataset that does not contain human or animal-related information. Therefore, ethical guidelines concerning human and animal studies are not applicable.
Declaration of Generative AI Use: We declare that no generative AI or AI-assisted technologies were used in the writing of this manuscript.